# Combinatorial metaplexes and centrality indices for identifying higher-order interactions

### v1

@@ Abstract
Simplicial complexes provide a combinatorial framework for modelling higher-order interactions (HoIs), that is, interactions involving more than two units. In this representation, simplices encode interactions and inclusion relations determine the structural organisation of the system. In classical graph-theoretic and simplicial complex-based models, vertices are treated as structurally homogeneous entities, and their intrinsic properties do not influence the formation of higher-dimensional simplices. However, in many real-world systems, HoIs are influenced by discrete intrinsic properties of the participating units. To address this limitation, we introduce the combinatorial metaplex, consisting of two interacting components: an underlying simplicial complex serving as an admissibility structure that specifies boundary-compatible higher-order simplex candidates, and a concentration layer that assigns concentration to vertices, extends it to simplices, and induces a threshold-based rule governing the realisation and weighting of higher-dimensional simplices. We investigate how simplex concentration influences the emergence of HoIs and identify the true HoIs as those satisfying both combinatorial admissibility and concentration-based realisation. Using facet-mediated adjacency, we define degree and closeness centrality for non-facet simplices. We further introduce a one-parameter family of centralities that interpolates between purely structural connectivity and concentration-induced coupling. This provides a framework for analysing structurally enriched higher-order networks.
@@ 1  Introduction
Systems characterised by multiple interacting components and collective phenomena, including nonlinearity, emergence, and adaptive behaviour, are central to complex systems theory [25, 35, 52, 17]. These systems are often modelled using graph-theoretic representations, leading to complex networks in which system units are represented as vertices and their interactions as edges [32, 23, 47, 10, 14]. Examples include protein–protein interaction networks, brain networks, and ecological networks [32].
@@
However, many complex systems exhibit interactions involving more than two units, in addition to pairwise interactions, referred to as higher-order interactions (HoIs) [4, 11, 16, 12]. To represent HoIs explicitly, more general combinatorial structures such as hypergraphs and simplicial complexes are commonly employed [39, 3, 46]. A hypergraph consists of a set of vertices together with a collection of subsets of these vertices, called hyperedges. Unlike a graph, where each edge contains exactly two vertices, a hyperedge may contain any number of vertices. A simplicial complex is a collection of subsets of a vertex set that is closed under inclusion: whenever a set belongs to the collection, all of its subsets also belong to it. The elements of a simplicial complex are called simplices, and a simplex containing $k+1$ vertices is called a $k$-simplex. In this representation, HoIs are modelled as simplices (such as triangles, tetrahedra, and higher-dimensional analogues), with the closure property ensuring that all lower-dimensional faces of an interaction are also present [30].
@@
In many existing studies, representations of systems—such as graphs, hypergraphs, and simplicial complexes—provide powerful tools for modelling the interaction structure of a system [16, 10, 11]. However, these formalisms primarily capture patterns of interaction and typically disregard intrinsic properties of the individual units forming the system. In practice, internal attributes of system components often influence how interactions are formed and sustained [31].
@@
The metaplex framework, introduced by Estrada and collaborators, unifies internal structure and interaction structure within a single representation [22, 31]. In this setting, each vertex carries an internal structure that supports a continuous dynamical process, while interactions between vertices are modelled by a network. This enables the study of diffusion processes evolving both within components and across their connections.
@@
In many real-world systems, however, intrinsic properties of units are discrete and quantifiable rather than continuous [7, 6, 13]. Vertices may carry measurable attributes such as the $h$-index or citation count of an author in a co-authorship network, the population or economic capacity of a city in a transportation network, or the abundance of a species in an ecological corridor network [45, 43, 49]. These heterogeneous quantities influence interaction patterns but are not generated by internal dynamical processes. Studies on node-weighted networks demonstrate that such intrinsic vertex weights affect structural measures including centrality and flow [45].
@@
Many higher-order network models employ abstract simplicial complexes, in which higher-dimensional simplices arise purely from combinatorial inclusion [52]. In particular, in clique-based constructions every complete subgraph is interpreted as a HoI, regardless of vertex-level attributes [35]. Consequently, HoIs emerge solely from pairwise closure, without incorporating the intrinsic concentration of the participating units.
@@
This motivates the introduction of the combinatorial metaplex. In this framework, each vertex is endowed with an intrinsic structure representing the internal organisation of the corresponding unit. A concentration is defined on this structure as a scalar quantity capturing its intrinsic capacity. These vertex-level concentrations constitute the only fundamental numerical quantities in the model.
@@
Unlike multilayer network models [15, 29], which represent different types of pairwise relations across distinct but interconnected network layers, the combinatorial metaplex introduces heterogeneity directly at the vertex level by assigning intrinsic structure to individual units.
@@
Within this vertex-driven setting, higher-dimensional simplices are not assigned intrinsic weights. Instead, their weights are induced from the concentrations of their constituent vertices through the simplicial incidence structure, implemented via the fractional weight distribution principle [9, 34]. Because a vertex may participate in multiple simplices, its concentration is distributed among incident simplices according to this incidence relation. Consequently, any numerical value associated with a simplex arises solely from aggregated vertex contributions, and no independent weight is assigned at the simplex level. This distinguishes the combinatorial metaplex from weighted simplicial complexes (see, e.g., [44, 19, 5]), where a non-negative real-valued function is assigned directly to simplices, thereby prescribing numerical strengths to higher-order interactions.
@@
In combinatorial metaplex, the realisation of a higher-order simplex is governed by a structural–quantitative condition. A simplex is realised only if the aggregated concentration of its boundary simplices exceeds a prescribed threshold, which acts as a global reference parameter. Thus, HoIs arise from two simultaneous requirements: combinatorial admissibility and sufficient inherited concentration. This ensures that realised HoIs are supported by intrinsic vertex-level properties rather than appearing solely through combinatorial closure.
@@
Finally, we introduce centrality measures on combinatorial metaplexes — including degree, closeness, and harmonic centrality — in order to analyse how vertex-level intrinsic properties influence higher-order structural organisation.
@@
A simplicial complex of dimension $1$ (i.e., the $1$-skeleton of $\Delta$) is a graph, and various centrality indices for pairwise networks have been introduced in the literature [48, 47, 33, 40]. The most basic centrality, degree centrality, counts the number of incident edges (that is, edges containing a vertex) to a vertex. In an edge-weighted graph, the strength of a vertex is defined as the aggregate weight of its incident edges. In classical graph-theoretic formulations, centrality measures have primarily focused on the edge-based adjacency structure. The incorporation of vertex weights into centrality definitions has been explored in [1, 45, 2, 28]. Our work extends these constructions by defining centrality measures for fully weighted graphs (with both vertex and edge weights), thereby providing a coherent treatment of intrinsic and adjacency-based properties.
@@
Several notions of adjacency between simplices have been proposed in the literature. Centrality in higher-order networks provides a quantitative framework for identifying the most influential simplices (or vertices). Extending graph-based centralities to simplicial complexes requires an appropriate notion of adjacency between simplices. The adjacency between two distinct $k$-simplices has been defined in terms of upper adjacency (when there exists a $(k+1)$-simplex containing both) and lower adjacency (when there exists a (k1)(k-1)-simplex contained in both) [26]. A generalised notion of adjacency, in which two distinct $k$-simplices are lower adjacent but not upper adjacent, was introduced in [24]. The adjacency between arbitrary simplices were defined in [41, 42].
@@
The generalised degree centrality in $d$-exclusive simplicial complex of $\Delta$ (i.e.,family of $d$-simplices of $\Delta$ and their faces) was defined in [20] and subsequently extended to betweenness and closeness centralities in [36]. Using a tensorial representation of the $d$-exclusive simplicial complex, the strength of a simplex in a weighted simplicial complex was defined in [21]. In defining centralities for the combinatorial metaplex, we follow the facet-based adjacency framework proposed in [37], in which two distinct $k$-simplices are said to be adjacent whenever they are contained in a common facet. Characterising simplices with maximal centrality yields structural insight into the combinatorial metaplex and consequently, into the underlying system it models.
@@
The main contributions of this work are as follows: (1) We introduce the combinatorial metaplex, a discrete two-layered construction that couples intrinsic vertex-level data with higher-order simplicial structure. (2) We extend the vertex concentration map to the entire simplicial complex via a fractional contribution map, propagating intrinsic vertex properties to higher-dimensional simplices through the incidence structure. (3) We establish a conservation relation between vertex-level concentrations and maximal higher-order simplices: the total induced weight of all facets coincides exactly with the total concentration assigned to the 0-simplices. (4) We develop a concentration-based framework for reconstructing simplicial complexes from data. Unlike clique-based constructions, which treat every clique of size greater than two as a HoI, our approach identifies only those simplices whose aggregated boundary concentration exceeds a prescribed threshold, thereby detecting concentration-supported HoIs. (5) We define degree centrality for fully weighted graphs consistent with previous definitions in weighted graphs. (6) We introduce degree, closeness, and harmonic centralities for the combinatorial metaplex using a facet-mediated notion of adjacency and weighted walks.
@@ 3.1  The combinatorial metaplex
As mentioned previously, the metaplex framework introduced in [22], models the internal vertex structure using locally compact metric–measure spaces. These spaces naturally support continuous dynamics, such as diffusion. However, in many systems, the internal structure of units is inherently discrete and does not admit a meaningful continuous description. In such cases, replacing internal metric–measure spaces by a finite combinatorial structure provides a natural generalisation.
When vertices possess discrete intrinsic properties that affect the global topology of a system, and these properties can be expressed as a finite discrete combinatorial structure, we introduce the discrete metaplex framework. Within this framework, each vertex is associated with an internal discrete structure endowed with a discrete measure, which we formalise by defining a concentration map on the set of vertices.
@@
Given a CM as in (1), for a vertex $v_{j}\in V,\ 0\leq i\leq n$, let $\mathcal{I}(v_{j})=Y_{j}\in\mathcal{Y}$. The concentration is a map
@@
In the CM framework, vertices are endowed with a combinatorial structure that encodes their intrinsic discrete properties. This internal structure is represented by strictly positive rational values. By assigning positive weights, we ensure that the contribution of each vertex is strictly positive and is inherited by higher-dimensional simplices in the hierarchical extension.
Henceforth, the term ‘weight’ will also be used to denote concentration. In the subsequent sections, we develop a method to extend the weight map from the vertex set to the simplicial complex.
@@ 3.2  The weight map on 0-simplices
For convenience, we again denote the map $a^{\prime}$ by $a$, unless this causes ambiguity in notation. With the vertices now distinguished by their concentration values, we next define how this weight extends to higher-dimensional simplices (edges, triangles, etc.), thereby yielding a fully weighted simplicial structure.
@@ 3.3  The fractional weight contribution map and contribution numbers
For convenience, we also denote the extended map on $S_{1}$ again by $a$, unless this causes ambiguity in notation. Finally, the weight of a $1$-simplex $\sigma^{(1)}\in S_{1}$ is defined as the sum of the contributions of its constituent vertices. Thus, extending the weight map $a$ to $S_{1}$, we obtain
@@ 3.6  Comparison of the total weights of 0-simplices and facets
Substituting this expression into the left-hand side yields
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})=\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)})-\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-1\end{subarray}}a(\sigma)$ (4)
Rearranging the terms gives,
$\sum_{i=1}^{n_{q}}a(\sigma_{i}^{(q)})+\sum_{\begin{subarray}{c}\sigma\text{ is a facet of }\Delta\\ dim(\sigma)=q-1\end{subarray}}a(\sigma)=\sum_{i=1}^{n_{q-1}}a(\sigma_{i}^{(q-1)}),$
which completes the proof. ∎
@@ 3.7  Inference of higher-order simplices
Within the CM framework, we introduce a boundary-based deterministic rule for admitting higher-order simplices into the complex. For $q\geq 2$, the inclusion of a combinatorially feasible $q$-simplex $\sigma^{(q)}$ is determined solely from the weights of its boundary (q1)(q-1)-simplices.
@@
This defines a concentration-based CM model in which higher-order simplices are added through a deterministic threshold criterion.
In the clique complex construction, the existence of a $q$-simplex is determined purely combinatorially: a simplex is present if and only if all of its (q1)(q-1)-faces are present. Equivalently, the vertices of the simplex span a clique in the $1$-skeleton.
In the CM, this combinatorial feasibility condition is retained but refined by weight information. While the presence of all boundary (q1)(q-1)-simplices guarantees clique compatibility, an additional structural constraint is imposed: a $q$-simplex is admitted only if its aggregated boundary weight exceeds the prescribed threshold $\theta_{q}$.
Consequently, the deterministic threshold criterion filters purely combinatorial closure, selecting only those simplices that exhibit sufficient collective boundary support.

### v2

@@ Abstract
Complex systems consist of interacting units whose interactions may be pairwise, involving two units, or higher-order, involving more than two units simultaneously. Graphs capture pairwise interactions and represent such systems as networks, whereas simplicial complexes can capture higher-order interactions (HoIs) and represent them as higher-order networks comprising simplices. In the clique complex construction, HoIs arise whenever vertices form a clique in the underlying graph. In classical graph-theoretic and simplicial-complex models, vertices are treated as structurally indistinguishable objects. However, in many real-world systems vertices possess internal structure, and their intrinsic properties influence the HoIs present in the system. To address this limitation, we introduce the combinatorial metaplex, consisting of two interacting components: an underlying simplicial complex that serves as an admissibility structure specifying boundary-compatible higher-order simplex candidates, and a concentration layer defined by a concentration map assigning a value to each vertex and extending the map to simplices so that the resulting distribution satisfies a conservation relation between vertex weights and facet weights. This concentration layer provides a deterministic threshold rule governing the inclusion of true HoIs. Using facet-mediated adjacency and weighted walks, we define one-parameter families of degree, closeness, and harmonic centralities for non-facet simplices, interpolating between those determined solely by the simplicial complex and those determined by concentration-induced coupling. The framework is illustrated through a representative example, including a comparison of HoIs obtained from the clique complex and the combinatorial metaplex, followed by an analysis of edge centralities within the combinatorial metaplex.
@@ 1  Introduction
In the study of complex systems, which are characterised by many interacting units, understanding the structure of interactions among these units is fundamental. Such systems often exhibit collective phenomena such as nonlinearity, emergence, and adaptive behaviour [25, 38, 54, 16]. When interactions occur pairwise, these systems are often modelled using graph-theoretic representations, leading to complex networks in which units are represented as vertices and their interactions as edges [35, 23, 50, 10, 13]. Examples of such networks include protein–protein interaction networks, brain networks, and ecological networks [35].
@@
However, many complex systems exhibit interactions among the units of group size greater than two, in addition to pairwise interactions. Such interactions are referred to as higher-order interactions (HoIs) [4, 11, 15, 12]. To represent HoIs explicitly, more general combinatorial structures such as hypergraphs and simplicial complexes are commonly employed [42, 3, 49]. A hypergraph consists of a finite set of vertices together with a collection of subsets of these vertices, called hyperedges. Unlike a graph, where each edge contains exactly two vertices, a hyperedge may contain any number of vertices. A simplicial complex is a collection of subsets of a vertex set that is closed under inclusion: whenever a set belongs to the collection, all of its subsets also belong to it. The elements of a simplicial complex are called simplices, and a simplex containing $k+1$ vertices is called a $k$-simplex. In this representation, HoIs are modelled as simplices (such as triangles, tetrahedra, and higher-dimensional analogues), with the closure property ensuring that all lower-dimensional faces of an interaction are also present [32].
@@
In many existing studies, mathematical representations such as graphs, hypergraphs, and simplicial complexes are widely used to model the interaction structure of complex systems [15, 10, 11]. However, these models primarily describe the structure of interactions and disregard intrinsic properties of the individual units forming the system. In many real-world systems, the internal properties of units affect the formation of HoIs [33].
@@
The metaplex framework, introduced by Estrada and collaborators, unifies internal structure and interaction structure within a single representation [22, 33]. In this setting, each vertex carries an internal structure that supports a continuous dynamical process, while interactions between vertices are modelled by a network. This enables the study of diffusion processes that evolve both within components and across their connections.
@@
In some systems, however, intrinsic properties of units are discrete and quantifiable rather than continuous [7, 6, 34, 19]. Vertices may carry measurable attributes such as the $h$-index or citation count of an author in a co-authorship network, the population of a city in a transportation network, or the abundance of a species in an ecological corridor network [48, 46, 19]. These heterogeneous quantities influence interaction structure but are not generated by internal dynamical processes [33]. Studies on vertex-weighted networks demonstrate that such intrinsic vertex weights affect structural measures including centrality and flow [48].
@@
Many higher-order network models employ abstract simplicial complexes, in which higher-dimensional simplices arise purely from combinatorial inclusion [54]. In particular, in clique-based constructions every complete subgraph is interpreted as a HoI, regardless of vertex attributes [38, 26]. Consequently, HoIs emerge solely from pairwise closure, without incorporating the intrinsic properties of the participating units.
@@
This motivates the introduction of the combinatorial metaplex. In this framework, each vertex is endowed with a combinatorial structure representing the internal structure of the corresponding unit. A concentration is assigned to each vertex as a scalar quantity representing an intrinsic quantitative property of the unit. These vertex concentrations constitute the fundamental numerical quantities in the model.
@@
Unlike multilayer network models [14, 31], which extend networks outward through multiple layers and interlayer edges connecting vertices across these layers, the combinatorial metaplex introduces vertex-level heterogeneity by assigning internal structure to individual units.
@@
Within this vertex-driven setting, higher-dimensional simplices are not assigned primary weights. Instead, their weights are induced from the concentrations of their constituent vertices through the simplicial inclusion structure, implemented via the fractional weight distribution principle [9, 37]. Because a vertex may participate in multiple simplices, its concentration is distributed among incident simplices (that is, simplices containing that vertex) according to this inclusion relation. Consequently, any numerical value associated with a simplex arises solely from aggregated vertex contributions, and no independent weight is assigned at the simplex level. This distinguishes the combinatorial metaplex from weighted simplicial complexes (see, e.g., [47, 18, 5]), where a non-negative real-value is assigned directly to simplices, thereby prescribing numerical strengths to HoIs.
@@
In the combinatorial metaplex, the inclusion of a higher-order simplex is governed by a structural–quantitative condition. A simplex is included only if the aggregated concentration of its boundary simplices exceeds a prescribed threshold. Thus, simplices arise from two simultaneous requirements: combinatorial admissibility and sufficient inherited concentration. This ensures that realised true HoIs are supported by intrinsic vertex-level properties rather than appearing solely through combinatorial closure.
@@
Finally, we introduce centrality measures on combinatorial metaplexes—including degree, closeness, and harmonic centralities—to analyse how intrinsic vertex-level properties influence the structure of HoIs.
@@
A simplicial complex of dimension $1$ (or the $1$-skeleton of a simplicial complex $\Delta$) is a graph, and various centrality indices for pairwise networks have been introduced in the literature [51, 50, 36, 43]. The most basic centrality, degree centrality, counts the number of incident edges to a vertex (that is, edges containing that vertex). In an edge-weighted graph, the strength of a vertex is defined as the aggregate weight of its incident edges. In classical graph-theoretic formulations, centrality measures have primarily focused on the weighted edge-based adjacency structure. The incorporation of vertex weights into centrality definitions has been explored in [1, 48, 2, 30]. Our work extends these constructions by defining centrality measures for fully weighted graphs, in which both vertices and edges are assigned weights, thereby providing a framework that couples intrinsic vertex properties with adjacency-based network structure.
@@
Centrality measures in higher-order networks quantify the importance of simplices within the interaction structure. Extending graph-based centralities to simplicial complexes requires an appropriate notion of adjacency between simplices. The adjacency between two distinct $k$-simplices has been defined in terms of upper adjacency (when there exists a $(k+1)$-simplex containing both) and lower adjacency (when there exists a (k1)(k-1)-simplex contained in both) [27]. A generalised notion of adjacency, in which two distinct $k$-simplices are lower adjacent but not upper adjacent, was introduced in [24]. The adjacency between arbitrary simplices was defined in [44, 45].
@@
The generalised degree centrality in the $d$-exclusive simplicial complex of $\Delta$ (i.e., the family of $d$-simplices of $\Delta$ and their faces) was defined in [20] and subsequently extended to betweenness and closeness centralities in [39]. Using a tensorial representation of the $d$-exclusive simplicial complex, the strength of a simplex in a weighted simplicial complex was defined in [21]. In defining centralities for the combinatorial metaplex, we adopt the facet-based definition of adjacency proposed in [40], in which two distinct $k$-simplices are said to be adjacent whenever they are contained in a common facet.
@@
The main contributions of this work are as follows: (1) We introduce the combinatorial metaplex, a two-layered construction that couples intrinsic vertex-level data with higher-order simplicial structure. (2) We extend the vertex concentration map to the entire simplicial complex via a fractional contribution map, propagating intrinsic vertex properties to higher-dimensional simplices through the incidence structure. (3) We establish a conservation relation between vertex concentrations and maximal higher-order simplices: the total induced weight of all facets equals the total concentration assigned to the 0-simplices. (4) We develop a concentration-based framework for constructing simplicial complexes from pairwise data. Unlike clique-based constructions, which consider every clique of size greater than two as a HoI, our approach identifies only those simplices whose aggregated boundary concentration exceeds a prescribed threshold, thereby detecting concentration-supported HoIs. (5) We define degree centrality for fully weighted graphs consistent with previous definitions in weighted graphs. (6) We introduce degree, closeness, and harmonic centralities for the combinatorial metaplex using a facet-mediated notion of adjacency and weighted walks. (7) We illustrate the proposed framework using a real-world network example. This example demonstrates the construction of the combinatorial metaplex and shows how the concentration-based inclusion rule identifies true HoIs, in contrast to the purely combinatorial HoIs obtained from the clique complex.
@@ 3.1  The combinatorial metaplex
As mentioned previously, the metaplex framework introduced in [22] models the internal vertex structure using locally compact metric–measure spaces. These spaces naturally support continuous dynamics, such as diffusion. However, in many systems, the internal structure of units is inherently discrete and does not admit a meaningful continuous description. In such cases, replacing internal metric–measure spaces with a finite combinatorial structure provides a natural generalisation. To model such systems, we introduce the combinatorial metaplex framework. In this framework, each vertex is associated with a finite combinatorial internal structure endowed with a discrete measure, which is formalised through a concentration map defined on the set of vertices.
@@
Given a CM as in (1), for a vertex $v_{j}\in V,\ 0\leq j\leq n$, let $\mathcal{I}(v_{j})=Y_{j}\in\mathcal{Y}$. The concentration is a map
@@
In the CM framework, vertices are endowed with a combinatorial structure that encodes their intrinsic discrete properties. The concentration map assigns to each vertex the discrete measure of its internal structure, that is, the total weight of the elements in the associated set. Consequently, each vertex is assigned a strictly positive concentration and thus contributes non-trivially.
Henceforth, the term ‘weight’ will be used interchangeably with concentration. In the subsequent sections, we develop a method to extend the weight map from the vertex set to the simplicial complex.
@@ 3.2  The weight map on 0-simplices
For convenience, we denote the map $a^{\prime}$ again by $a$. With the vertices now distinguished by their concentration values, we next define how this weight extends to higher-dimensional simplices (edges, triangles, etc.), thereby yielding a fully weighted simplicial structure.
@@ 3.3  The fractional weight contribution map and contribution numbers
For convenience, we also denote the extended map on $S_{1}$ again by $a$. Finally, the weight of a $1$-simplex $\sigma^{(1)}\in S_{1}$ is defined as the sum of the contributions of its constituent vertices. Thus, extending the weight map $a$ to $S_{1}$, we obtain
@@ 3.6  Comparison of the total weights of 0-simplices and facets
Substituting this expression into the left-hand side and rearranging the terms gives
@@ 3.7  Inference of higher-order simplices
Within the CM framework, we introduce a boundary-based deterministic rule for admitting higher-order simplices into the complex. We assume that the realization of a higher-order simplex is supported by sufficiently large boundary weights. For $q\geq 2$, the inclusion of a combinatorially feasible $q$-simplex $\sigma^{(q)}$ is determined solely from the weights of its boundary (q1)(q-1)-simplices.
@@
This defines a weight-based CM model in which higher-order simplices are introduced through a deterministic threshold criterion.
In the clique complex construction, the existence of a $q$-simplex is determined purely combinatorially: a simplex is present if and only if the vertices of the simplex induce a complete subgraph of the $1$-skeleton of $\Delta$.
In the CM, this combinatorial condition is retained but refined using weight information. While the presence of all boundary (q1)(q-1)-simplices ensures that the simplex is combinatorially admissible, an additional structural constraint is imposed: a $q$-simplex is included only if its aggregated boundary weight exceeds the prescribed threshold $\theta_{q}$.
The threshold $\theta_{q}$ can be interpreted in terms of the average weight of (q1)(q-1)-simplices. The quantity $\bar{a}_{q-1}$ denotes the average weight of all (q1)(q-1)-simplices currently present in the simplicial complex. Since a $q$-simplex has exactly $(q+1)$ boundary (q1)(q-1)-faces, the value $(q+1)\bar{a}_{q-1}$ represents the total weight obtained when each boundary face carries the average weight. The threshold therefore serves as a reference level for evaluating the collective boundary support of a candidate simplex. A $q$-simplex is included only when its aggregated boundary weight exceeds this thresold.
Consequently, in the CM the deterministic threshold criterion filters out purely combinatorial closures and includes only those simplices whose boundary weight exceeds the prescribed threshold.